OpenStudy (wavyapples):

Owen receives $10$ and puts it into his savings account. He adds $0.50 to the account each day for a number of days, dd, after that. He writes the expression 10+0.5(d−1) to find the amount of money in his account after d days. Which statement about his expression is true?

OpenStudy (jadenwalker76):

well it depends on the number of days but here's how you do it. you use d as a substitute for the number of days so.... 10+0.5(d-1)= 10+0.5(1-1)=10 10+0.5(d-1)=10+0.5(2-1)=10.5 10+0.5(d-1)=10+0.5(3-1)=11 and you just keep on plugging the next number up in replace for d for the number of days that there are. I hope this helped..